OpenStudy (princesssleelee):

Given that f(x) = 2x + 1 and g(x) = 3X-1/2 solve for g(f(3)).

OpenStudy (er.mohd.amir):

f(3)=put x=3 in f(x) then f(3)=ans then put ans in g(x)=g(ans of f(3))= u ans

OpenStudy (princesssleelee):

That doesnt make much sense to me at all @Er.Mohd.AMIR

OpenStudy (er.mohd.amir):

f(x)=2(3)+1=6+1=7 now find g(7) put x=7 in g(x)

OpenStudy (princesssleelee):

So the answer would be seven

jhonyy9 (jhonyy9):

so for you can getting f(3) you need substituting the 3 in place of x and in case of f(x) and this value of f(3) you need substituting in place of x in case of g(x) ok ?

OpenStudy (princesssleelee):

This is just one form of math that I have completely no idea on how to do it...

OpenStudy (princesssleelee):

I got the answer 7. I think..

OpenStudy (princesssleelee):

@jhonyy9

OpenStudy (er.mohd.amir):

no ans is not 7 put 7 in place of x in g(x) and solve

OpenStudy (er.mohd.amir):

g(f(x))=g(7)=3*7-1/2=21-1/2=41/2=answer

jhonyy9 (jhonyy9):

so 7 is just f(3) =7 but you need to calculi g(f(3) = ? yes ?

OpenStudy (princesssleelee):

x=5

jhonyy9 (jhonyy9):

so because f(3)=7 so than you need substituting 7 in place of x in case of g(x) ok ?

OpenStudy (princesssleelee):

Okay, Whenever I substituted 7, I got x=5

jhonyy9 (jhonyy9):

3x - 1 g(x) = ------ 2 hence is right ?

OpenStudy (princesssleelee):

Yes

jhonyy9 (jhonyy9):

so this mean that 5 is correct answer - yes ?

OpenStudy (princesssleelee):

I believe so lol

jhonyy9 (jhonyy9):

so do you understand it now why is this way to solve this exercise ?

OpenStudy (princesssleelee):

Yes I understand it now, for the most part :) thank you very much

jhonyy9 (jhonyy9):

was my pleasure good luck